OpenStudy (anonymous):

If a power station generates electricity at 20A at 5000V and the transmission lines has a resistance of 20ohms, determine the power lost to heat during transmission.

OpenStudy (anonymous):

If the power output from the station is connected to step up than has turns ratio 1:10 from primary to secondary, determine the power lost in the transmission line due to joule heating.

OpenStudy (anonymous):

The Power generated by the power station is P = VI = 100000 Watts = 10^5 Watts So as the no. of turns ratio is 1:10 the Voltage will rise 10 times and the Current will fall 10 times , using the formula P=VI since V inversely proportional to I or current. The stepped up Voltage will be 5000*10 = 50000 Volts and the Current will be 20/10 = 2 A Since, P = VI = (I^2)*R = 80 Watts , therefore the amount of energy or heat lost is 80 Watts which also supports the fact that electricity transmission is more efficient in higher voltages.

OpenStudy (anonymous):

The ratio is about what ? Voltage ? If so : \[P_{Station}=U_{generator}.I_{generator}=P_{Output}=U_{output}.I_{output}\] \[P_{Output}=(U_{generator}*10).I_{output}\] \[I_{output}=I_{generator}/10\] \[P_{Joule Effect}=(R.I_{Ouput}).I_{Output}=(20/10^2)*I_{generator}^2=80\]